# Mixed Reality Scenic Live Streaming for Cultural Heritage: Visual Interactions in a Historic Landscape

### v2

@@ Abstract.
journalyear: 2025copyright: noneconference: Proceedings of the International Conference on Human-Engaged Computing; November 21–23, 2025; Singaporebooktitle: Proceedings of the International Conference on Human-Engaged Computing (ICHEC ’25), November 21–23, 2025, Singaporedoi: XXXXXXX.XXXXXXXisbn: XXX-X-XXXX-XXXX-X/25/11submissionid: 11ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Interface design prototyping
@@ 2.1. XR for Cultural Heritage Sites
$2.$ Scenic Live Streams as a Design Space
@@
Figure 1. Our design process for an MRSLS at cultural heritage sites.:
@@ 3.3.1. Camera View
Figure 2. The initial overlooking view (left) is preferred by plain SLS like SkylineWebcams and Live China. The new first-person view (right) better fits interactive MRSLS.:
@@
(a) A screenshot in the middle of the session. The plain scenery of regular scenic live streaming is enhanced by virtual and real-time interactive elements.:
(b) The lotuses on the lake, each controlled by one viewer. The lotus above is shining. The blue lotus below is dashing and colliding with the white one.:
(c) The boat carrying top scorers’ names after the verse game.:
(d) An oil-paper umbrella with hashtagged comment.:
(e) A fish jumping out of water.:
(f) Fireworks on the sky.:
@@ 3.4. Technical Implementation
Figure 4. The software architecture of our MRSLS prototype.: A diagram of the MRSLS system implementation.
@@ 5.3. The MR Design: Visual and Functional
Figure 5. Questionnaire responses on MR design:
@@ 6.1. Unique Affordances of MR for Scenic Live Streams
Figure 6. MRSLS as a heritage site interaction experience encompasses three aspects.:

### v3

@@ Abstract.
journalyear: 2025copyright: noneconference: Proceedings of the 2025 International Conference on Human-Engaged Computing; November 21–23, 2025; Singapore, Singaporebooktitle: Proceedings of the 2025 International Conference on Human-Engaged Computing (ICHEC 2025), November 21–23, 2025, Singapore, Singaporedoi: 10.1145/3786995.3786999isbn: 979-8-4007-2234-9/2025/11submissionid: 11ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Interface design prototyping
@@ 2.1. XR for Cultural Heritage Sites
2.2. Scenic Live Streams as a Design Space
@@
Figure 1. Our design process for an MRSLS at cultural heritage sites.
A multi-stage design process flow diagram segmented in three parts. The first part focuses on cultural relevance. It contains a formative interview on conceptual and cultural association with West Lake and Hangzhou City, which produces virtual elements. The second part focuses on geographical integration. It contains filtering the virtual elements by technical feasibility, visual occlusion, and interference with the scenery focus. The third part focuses on participatory engagement. It contains feature brainstorming that produces an initial prototype, and iterative design studies that produce a final prototype.
@@ 3.3.1. Camera View
Refer to caption An arrow connects two pictures. On the left is an overlooking view from one side of the West Lake. The building and trees are far away. On the right is a first-person view of one side of West Lake. The buildings and trees are larger and clearer. The view also captures a bridge and some pedestrians.
Figure 2. The initial overlooking view (left) is preferred by plain SLS like SkylineWebcams and Live China. The new first-person view (right) better fits interactive MRSLS.
@@
Refer to caption A moment of the West Lake scene. In addition, there are some virtual fireworks exploding in the sky. There are some virtual lotuses floating on the water with ripples and name tags. Some game UI is overlaid on the left. The instructions are overlaid on the top right.
(a) A screenshot in the middle of the session. The plain scenery of regular scenic live streaming is enhanced by virtual and real-time interactive elements.
Refer to caption Three lotuses in red, blue, and white, respectively, are floating on the lake. The blue one is shining.
(b) The lotuses on the lake, each controlled by one viewer. The lotus above is shining. The blue lotus below is dashing and colliding with the white one.
Refer to caption A boat carrying a board with the top three scorers’ names is floating on the lake.
(c) The boat carrying top scorers’ names after the verse game.
Refer to caption An oil-paper umbrella is floating in the sky. A short piece of text is above the umbrella.
(d) An oil-paper umbrella with hashtagged comment.
Refer to caption A fish with a name tag jumping out of water.
(e) A fish jumping out of water.
Refer to caption Some yellow fireworks exploding in the sky.
(f) Fireworks on the sky.
@@ 3.4. Technical Implementation
Refer to caption A diagram of the MRSLS system implementation. A video stream from the camera is fed to a Unity program, which contains a “real-time” part and a “one-time” part. The one-time part uses the footage to calculate camera configurations and 3D reconstruction. The real-time part renders (1) the footage with proper camera configurations, (2) materials, shaders, and textures for collision and occlusion thanks to the 3D reconstruction, and (3) the MR elements that depend on the materials, shaders, and textures. The combined scene of the entire real-time part is sent to live-streaming platforms. Viewers’ activities (comments and gifts) are sent to both the live-streaming platform and the Unity program to control the MR elements.
Figure 4. The software architecture of our MRSLS prototype. Icons in this figure are designed by Freepik, Smashicons, Icons8, and Eucalyp.
@@ 5.3. The MR Design: Visual and Functional
Figure 5. Questionnaire responses on MR design
A Likert scale plot of user feedback on MR design. Almost all participants agree that “The MRSLS content design is appropriate”, “The MR elements are suitable for the scene”, “The MR elements embody the scenic uniqueness of West Lake”, “The MR elements match the local culture”, and “The art style is appropriate”. Slightly more than half of the participants think that “The art style is realistic” and “I feel as if interacting with the physical environment of West Lake”. Almost all participants DISAGREE that “I feel as if I changed the physical environment of West Lake”.
@@ 6.1. Unique Affordances of MR for Scenic Live Streams
Refer to caption A Venn diagram with three elements. The first, “Cultural”, contains “Tangible” and “Intangible”. The second, “Participatory”, contains “Individual”, “Collective”, and “Experiential”. The third, “Authentic”, contains “Livingness”, “Spatiality”, and “Context”.
Figure 6. MRSLS as a heritage site interaction experience encompasses three aspects.